OpenStudy (ahsome):

A wooden box of bolts has a mass of 250kg and requires a horizontal force of 2100N to slide it along a horizontal wooden surface at a constant speed. It has a coefficient of friction of 0.86.

OpenStudy (ahsome):

If the box were to be kept moving constantly at twice this speed what force would be needed to maintain this constant velocity

OpenStudy (anonymous):

First, find the normal contact force or N - it is the force that acts opposite to the force applied to the floor and in this case it is its weight. N=weight=mg=250*10=2500N Secondly, for a body to move at constant speed, it means that both driving (Fd) and Resistive (R) forces are equal, meaning R=Fd The R=$N, where $ is the coeffient of friction and your Fd is 2100N

OpenStudy (ahsome):

So, FR = 2150?

OpenStudy (shamim):

Frictional force=coefficient of friction*weight=0.86*250*9.8=?

OpenStudy (ahsome):

2107N

OpenStudy (ahsome):

So, if we keeping increasing the speed, the force never changes?

OpenStudy (shamim):

2107 is greater than 2100N

OpenStudy (anonymous):

Ahsome, is your g 10 or 9.8?

OpenStudy (ahsome):

9.8

OpenStudy (ahsome):

Wait, your answer is 2107

OpenStudy (shamim):

2107-2100=7N makes the body acceleration.not keep constant velocity

OpenStudy (ahsome):

Ok. So if we keep increasing the speed, the force doesn't change?

OpenStudy (anonymous):

Unless we can consider as rounded to 2100N, because the question announces that that force causes a constant speed, dont you think @shamim?

OpenStudy (ahsome):

I would say so. Its probably because i use 9.8

OpenStudy (shamim):

Ya. Frictional force will b equak to 2100N

OpenStudy (shamim):

So g will b less than 9.8. Right?

OpenStudy (ahsome):

Isn't it a bit more? 9.81 something

OpenStudy (shamim):

g=9.81 makes frictional force more than 2107N

OpenStudy (anonymous):

Guys, I think they have rounded.

OpenStudy (shamim):

Ya rounded

OpenStudy (anonymous):

So, having that considered, I say we foollow the momentum's formula: Ft/m=v, meaning if F doubles, v doubles, hence to maintain double the constant speed, the force has to be 2100*2=4200N.

OpenStudy (shamim):

No @Holos

OpenStudy (shamim):

Constant velocity but the value is doubled . Needs 2100N applied force

OpenStudy (shamim):

Frictional force=2100N Applied force=2100N Both r ewual n oposit So cancelled each other

OpenStudy (ahsome):

Exactly, @shamim. The velocity is doubled, but the force needed to sustain stays the same, since friction doesn't increase as you increase your speed.

OpenStudy (ahsome):

Thanks @shamim and @Hoslos. You guys were really helpful :)

OpenStudy (shamim):

U r most welcome

OpenStudy (ahsome):

:). Would you guys be able to help me with another question?

OpenStudy (shamim):

Of course. But u should ask question in main page

OpenStudy (ahsome):

Yah

OpenStudy (anonymous):

Yes, whenever I can possibly explain. I am here.

OpenStudy (ahsome):

I'l make a new question. See you guys there :)

OpenStudy (shamim):

Anyway the formula will b F*t=mv-mu F*t/m=v-u @Hoslos

OpenStudy (shamim):

And this formula is for changing velocity. Right? @Hoslos

OpenStudy (anonymous):

Yes, but since the v is constant, guess we can use v as the same.

OpenStudy (shamim):

When v is constant u will write v-u=0

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

Let us go help Ahsome, shall we? Let us keep warming up!